OpenStudy (anonymous):

Which equation justifies why 27^1/3 is equivalant to the cube root of 27?

OpenStudy (anonymous):

A) 33 = 27 B) 27 1 3 = 3 C) (27 1 3 )3 = 27 D) (27 1 3 )3 = (33 ) 1 3

OpenStudy (anonymous):

sorry i didnt put the power symbols

OpenStudy (anonymous):

A) 3^3 = 27 B) 27(1/3)= 3 C)(27(1/3))^3 = 27 D)(27(1/3))^3 = (3^3 )1/3

OpenStudy (anonymous):

A has the only true statement